40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 20 18:03:45 2011, Paul, Update, SUS, ITMY and SRM oplev calibrations - measured and estimated 
    Reply  Tue Sep 20 19:00:49 2011, Paul, Update, SUS, ITMY and SRM oplev calibrations - measured and estimated 
    Reply  Wed Sep 21 00:37:01 2011, rana, Update, SUS, ITMY and SRM oplev calibrations - measured and estimated 
       Reply  Wed Sep 21 09:10:15 2011, Paul, Update, SUS, ITMY and SRM oplev calibrations - measured and estimated 
Message ID: 5496     Entry time: Wed Sep 21 09:10:15 2011     In reply to: 5494
Author: Paul 
Type: Update 
Category: SUS 
Subject: ITMY and SRM oplev calibrations - measured and estimated 

Quote:

I found that some of the Optical Lever Servos were ON today and injecting nonsense into the interferometer optics. I have set all of the gains = 0 to save us more headaches.

Please leave them OFF until we review the servo and noise characterization results in the elog.

 I had previously set the gains to zero, see the first line of my entry on Monday 5468. I should have the servo and noise characterisation done today for these oplevs today, so we can review it soon.

ELOG V3.1.3-